# Bounding the conditional von-Neumann entropy for device independent cryptography and randomness extraction

### v1

@@ Abstract
This paper introduces a numerical framework for establishing lower bounds on the conditional von-Neumann entropy in device-independent quantum cryptography and randomness extraction scenarios. Leveraging a hierarchy of semidefinite programs derived from the Navascués-Pironio-Acín (NPA) hierarchy, our tool enables efficient computation of entropy bounds based solely on observed statistics, assuming the validity of quantum mechanics. The method’s computational efficiency is ensured by its reliance on projective operators within the non-commutative polynomial optimization problem. The method facilitates provable bounds for extractable randomness in noisy scenarios and aligns with modern entropy accumulation theorems. Consequently, the framework offers an adaptable tool for practical quantum cryptographic protocols, expanding secure communication possibilities in untrusted environments.
@@ I Introduction
In the realm of quantum mechanics, randomness is not merely a consequence of incomplete knowledge but an intrinsic property of nature itself, a principle famously highlighted by Max Born. This inherent randomness has profound implications for quantum information theory, particularly in applications that require the generation of secure random numbers or private cryptographic keys [1, 2]. The quantification of such randomness hinges on the concept of entropy, especially the conditional von-Neumann entropy:
$H(A|Q_{E})_{\rho_{AQ_{E}}}\coloneqq S(\rho_{AQ_{E}})-S(\rho_{Q_{E}}).$ (1)
Here, $\rho_{AQ_{E}}$ represents a classical-quantum state in the post-measurement system $A$ on Alice’s side, with quantum side-information $Q_{E}$, which we assume is fully accessible to an eavesdropper. The conditional von-Neumann entropy measures the maximum randomness extractable in a secure, asymptotic manner from quantum systems under certain constraints. Given its importance, lower bounding the conditional von-Neumann entropy has become a central task in both theoretical studies and practical applications of quantum information and especially security proofs in quantum cryptography [3, 4, 5].
@@
Recent advancements in quantum cryptography, particularly in device-independent quantum key distribution (DIQKD) and randomness extraction [6, 7, 8, 9, 10], harness the intrinsic unpredictability of quantum systems to secure communication against eavesdropping adversaries. DIQKD protocols rely on the principle that correlations observed in measurement outcomes from entangled quantum systems cannot be accounted for by any local hidden variable theory, as shown by violations of Bell inequalities [11, 12, 13, 14]. This fundamental characteristic implies that a conditional distribution
$p(a,b|x,y)$ (2)
@@
cannot be reproduced if $a$ and $b$ were deterministic outputs in hands of an evesdropper for each input pair $(x,y)$, with values governed by a shared random variable, like a common coin, which randomizes $a$ and $b$ jointly [15]. Instead, quantum non-locality, revealed through such correlations, ensures that outcomes cannot be explained by classical models, thus establishing true randomness by laws of quantum theory. The DIQKD approach offers an additional layer of security by requiring no trust in the measurement devices themselves. Given this framework, device-independent protocols make the accurate bounding of entropy under non-asymptotic conditions a central challenge [16, 17, 18]. The need to rigorously quantify the randomness of measurement outcomes, especially in real-world scenarios where noise and finite sample sizes introduce deviations, underscores the importance of non-asymptotic entropy bounds. Such bounds provide critical insights into the security achievable in device-independent settings, where only the observed distribution and the validity of quantum theory are assumed [19, 20, 21, 22, 23]. As research advances, these entropy bounds are becoming essential for verifying and implementing DIQKD protocols in practical cryptographic systems, thus expanding the possibilities of secure communication in untrusted environments [24, 25, 26].
@@
In this work, we contribute to the field of device-independent randomness extraction, addressing how much usable randomness can be extracted from a given experiment. We introduce a new numerical method that provides provable lower bounds for the extractable randomness in (1). Thus, we answer the question of how much randomness can be extracted from a specific type of experiment, assuming only the availability of measurement results from Alice and Bob and the premise that the experiment is governed by quantum theory. Our method complements the existing work of Brown, Fawzi, and Fawzi [3], as well as Tan et al. [4]. The demand for such tools has increased recently due to significant progress in entropy accumulation theorems [19, 20, 21, 22, 23], which provide concrete solutions for randomness extraction in finite-size regimes by reducing it to the asymptotic quantity in (1). Thus, after applying an entropy accumulation theorem, the remaining task is to calculate provable lower bounds for the conditional von-Neumann entropy (1) for specific noise models and input-output data scenarios. Furthermore, by confining the analysis to projective operators, our method achieves a high degree of computational efficiency, making it suitable for real-world applications.
@@ II Bounding the von-Neumann entropy
$D(\rho\|\sigma)\coloneqq\begin{cases}\operatorname{tr}[\rho\log\rho-\rho\log\sigma]&\ker[\sigma]\subseteq\ker[\rho]\\ +\infty&\text{else}\end{cases}$ (3)
by P. E. Frenkel [27] for two arbitrary positive operators $\sigma,\rho\in\mathcal{P}(\mathcal{H})$ on a Hilbert space $\mathcal{H}$
@@
For the rest of the work we use the expression $\operatorname{tr}^{-}[A]$ for the trace of the negative part of a self-adjoint operator $A$ on $\mathcal{H}$ with unique Jordan decomposition $A=A_{+}-A_{-}$ and positive operators $A_{+}$ respectively $A_{-}$. Similarly we define $\operatorname{tr}^{+}[A]$. As shown by Jenčová [28], the formula in (4) is even for infinite-dimensional separable Hilbert spaces true and can be rewritten in the following simple form
@@
$+\ln\lambda+1-\lambda\big{)}.$
The integral bounds $\mu,\lambda\in\mathbb{R}_{\geq 0}$ are solutions for the operator valued inequality
@@
The value of $\lambda$ can be found by solving the equation $\log\lambda=D_{\operatorname{max}}(\rho\|\sigma)$, which defines the max-relative entropy, $D_{\operatorname{max}}$ [29]. For the remainder of the work, we assume a tripartite experiment involving Alice, Bob, and Eve, modeled by the axioms of quantum theory as a tripartite state, a positive operator on a Hilbert space with trace equal to $1$,
@@
Importantly, the argument for the device independence specifies nothing concrete about the Hilbert space $\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{QE}$. In the following we follow the example of generating a raw-key in a DIQKD protocol, even though our methods can be applied for a larger class of problems section III. Afterwards, we discuss briefly some simple generalizations.
The fact that we consider the von-Neumann entropy means that we do not have to deal with finite size effects. In section III we argue this assumption for certain QKD scenarios. In a DIQKD protocol we usually have some measurements $\mathcal{X}$ on $\mathcal{S}(\mathcal{H}_{Q_{A}})$ with finite outcome set $A$ and similarly $\mathcal{Y}$ on $\mathcal{S}(\mathcal{H}_{Q_{B}})$ with finite outcome set $B$ respectively. Measurements are generally understood as sets of POVMs $\{M_{a|x}\}_{a\in A}$ for each $x\in\mathcal{X}$ and similarly $\{M_{b|y}\}_{b\in B}$ for each $y\in\mathcal{Y}$. In a more abstract language we understand it as a quantum channel from a commutative algebra into our observable algebra. In the most simplest form of a protocol, the raw-key is now constructed by one specific measurement $\tilde{x}\in\mathcal{X}$. This can be generalized later on [30], but we skip the technical details of it for now. In a next step, we use the fact that the asymptotically extractable randomness out of the fixed measurement $\tilde{x}\in\mathcal{X}$ can be described as [^2] For simplicity we abbreviate $\rho_{AQ_{E}|\tilde{x}}\equiv\rho_{AQ_{E}|X=\tilde{x}}$
@@
This yields that we are particular interested in lower bounding $-D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$. Specifically, device independent lower bounds satisfy the following prototypical type of optimization problem
@@
whereby we do not aim to specify the concrete underlying Hilbert space. Moreover, the second line can be understood as information about the joint statistics of Alice and Bob, which is nothing more than (2). A simple but sufficient and common example would be that the $c_{abxyi}\in\mathbb{R}$ coefficients form a Bell inequality [31, 3, 18]. In the following we need to estimate the program (9) with an actually quantifiable program. For this purpose, we use the integral representation (5) and derive an upper bound for the relative entropy (3) with an operator valued optimization problem.
@@
$+\ln\lambda+1-\lambda).$
@@
The Theorem II.1 provides a sequence of upper bounds that converge to the relative entropy. Specifically, this sequence of optimization problems can be understood as an application of [32, Appendix E]. We emphasize that a finite $\lambda$ is, at least in finite dimensions, equivalent to the statement that the relative entropy is finite. This is a crucial component in these optimization problems, allowing for effective numerical analysis. In the case of conditional von-Neumann entropy, a finite $\lambda$ always exists, as for any bipartite state $\rho_{AB}$ we have the fundamental operator inequality (see, e.g., [33, Appendix A]):
@@
Now, let us examine the specific optimization problem in (9). It can be lower-bounded using Theorem II.1 as an optimization over operators $P_{0},\ldots,P_{r}$. In particular, we observe that the optimal operators $P_{k}$, $0\leq k\leq r$, are projections, as they correspond to projections onto the positive part in the Jordan decomposition of the Hermitian operator $\alpha_{k}\rho+\beta_{k}\sigma$. Furthermore, by substituting the specific operators $\rho=\mathds{1}_{A}\otimes\rho_{Q_{E}}$ and $\sigma=\rho_{AQ_{E}|\tilde{x}}$, we find that $\lambda=1$ and $\mu$ can be set to 00 without loss of generality (see Appendix A for details). Thus, we can rewrite (9) as
@@
An application of Naimark’s dilation theorem (e.g., [34, Chap. Dilation Theory]) additionally allows us to rewrite the state $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$ without loss of generality as a pure state $\psi_{Q_{A}Q_{B}Q_{E}}\in\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}}$, with all measurements also being projective. By carefully examining the classical-quantum structure of $\rho_{AQ_{E}|\tilde{x}}$, the post-measurement state in our specific key basis $\tilde{x}\in\mathcal{X}$ can then be used to reformulate (14) as a non-commutative polynomial optimization in Theorem II.2.
@@
The polynomial optimization problem (15) can be interpreted as an optimization over the algebraic constraints within it. In a finite-dimensional Hilbert space setting, these constraints can be viewed as generating rules of a $\star$-subalgebra of $\mathcal{B}(\mathcal{H})$, where $\mathcal{H}$ is a Hilbert space (in infinite dimensions, caution is required to handle the topological closure of the algebra, though this remains feasible). Assuming that these relations form an algebra $\mathcal{A}$, we employ the Navascues-Pironio-Acin (NPA) hierarchy [35], applied in the algebraic framework as demonstrated in [36], to compute outer approximations to the optimization problem (15), thus lower bounds.
The approach involves relaxing the positivity requirement for linear, continuous functionals on $\mathcal{A}$. Rather than requiring positivity across the positive cone $\mathcal{A}_{+}$, we restrict it to a sum-of-squares cone $\Sigma_{2}^{(n)}\subset\mathcal{A}_{+}$ at level $n$. Optimization over this relaxed SDP can then be viewed as optimizing the dual cone $(\Sigma_{2}^{(n)})^{\star}$, which is representable in SDP form and provides a convergent hierarchy of outer approximations to (15). For further details, we refer to [37, 35, 36, 38] and Appendix D. In our applications, we use the software package from [39].
0.750.750.750.750.80.80.80.80.850.850.850.85000.50.50.50.51111CHSH valueconditional entropylower boundanalytic bound
0.750.750.750.750.80.80.80.80.850.850.850.85000.50.50.50.511111.51.51.51.5CHSH valueconditional entropylower bound $2$
0.80.80.80.80.90.90.90.911110011112222white noiseconditional entropylower bound 2322
Figure 1: The plots illustrate examples of lower bounds for conditional entropies across different scenarios. In the leftmost plot, we present a proof-of-principle example for lower bounds on the conditional entropy $H(A|X=0,Q_{E})_{\rho}$, depending on various values of the CHSH violation in a $2$-scenario (where each of two parties can choose between two measurements with two outcomes). This example is particularly significant because it allows for a verification of the general correctness of the method; as shown in [9], there are analytical values available for this scenario, enabling comparisons with computed results to validate accuracy. The middle plot displays the lower bounds on global randomness, specifically the conditional entropy $H(AB|X=0,Y=0,Q_{E})$, again in a $2$-scenario. Here, the dependency is on the CHSH inequality, and this plot serves to examine the impact of the CHSH violation on the joint conditional entropy of outcomes $A$ and $B$ given measurement settings $X=0$ and $Y=0$. Finally, the rightmost plot demonstrates the quantity $H(AB|X=0,Y=0,Q_{E})$ in a 2322-scenario, where each party has two possible outcomes and Bob has three measurement choices. In this case, we condition on the full distribution of a Werner state depending on dephasing noise.

### v2

# Reliable Entropy Estimation from Observed Statistics for Device-Independent Quantum Cryptography

@@ Abstract
We introduce a numerical framework for reliably estimating conditional von Neumann entropies, a central quantity in device-independent quantum cryptography and randomness generation. Our method is based on semidefinite relaxations derived from the Navascués–Pironio–Acín hierarchy and provides entropy bounds directly from observed statistics, assuming only the validity of quantum mechanics. The approach builds on a recent integral representation of entropy and substantially improves efficiency: it requires about half as many support variables as existing methods, while allowing these variables to be chosen projectively, leading to a significant reduction in runtime. These advances enable the practical certification of randomness and security even in noisy conditions, while integrating seamlessly with modern entropy accumulation theorems. Consequently, our framework becomes a versatile tool for quantum cryptographic protocols, broadening the possibilities for secure communication in untrusted environments.
@@ I Introduction
Almost a century ago, Max Born [1] famously suggested that the unpredictability observed in individual quantum measurements cannot always be explained by an incomplete knowledge about the experimental setup alone. Instead, this unpredictability can be an intrinsic property of nature itself—that is, true randomness exists. This idea is strikingly demonstrated by a successful Bell test.
Regarded from the perspective of quantum mechanics, a maximal violation of a Bell inequality like CHSH [2] demonstrates two key points. On one hand, it rules out the existence of hidden variables that could secretly determine a outcome. On the other, it shows that the correlated yet individually unpredictable measurement outcomes—emerging from a pure, maximally entangled state—are truly random and uncorrelated with any third party not involved in the experiment.
Today, Bell-type correlation experiments that are capable of closing a wide range of loopholes have moved beyond mere thought experiments—they are now a practical reality (see [3, 4, 5, 6, 7, 8] and references therein). However, since all physical experiments must contend with noise and decoherence, a maximal violation indicative of a pure state will most likely remain an idealized model. This necessitates the task of characterizing and certifying the amount of randomness present in observed noisy real-world data—an issue we address in this work.
Recognizing that the outcomes of a quantum measurement can be genuinely random marked not only a fundamental shift in our understanding of the world but also, much later, paved the way for the development of numerous information-theoretic protocols that would be impossible to achieve under the framework of any classical theory [9]. Most prominently, the field of quantum cryptography evolved from this, which by today ranges from the development of near-commercial technologies to foundational research pushing the limits of current capabilities.
Figure 1: Basic setting of a quantum mechanical correlation experiment for the generation of randomness or secret key. A source distributes a quantum state $\rho_{Q_{A}Q_{B}Q_{E}}$ between two legit users Alice and Bob, and an adversary Eve. Depending on inputs $x$ and $y$, Alice and Bob perform measurements resulting in classical random variables AA and $B$ to which Eve holds quantum side information. Our main focus is to estimate the securely asymptotically extractable randomness of AA and/or $B$ based on only considering the observed probability distributions $p(a,b|x,y)$.
@@
Recent advancements in quantum cryptography, particularly in device-independent quantum key distribution (DIQKD) and randomness extraction (see e.g. [10, 11] and references therein), harness the intrinsic unpredictability of quantum systems to secure communication against eavesdroppers. DIQKD protocols rely on the principle that correlations observed in measurement outcomes from entangled quantum systems cannot be explained by any local hidden variable theory, as shown by violations of Bell inequalities.
This fundamental characteristic implies that the conditional distribution
$p(a,b|x,y),$ (1)
@@
see Fig. 1, cannot be reproduced if $a$ and $b$ were deterministic outputs available to an eavesdropper for each input pair $(x,y)$, with values governed by a shared random variable, such as a shared coin toss. Instead, quantum non-locality—revealed through such correlations—ensures that outcomes cannot be explained by classical models. This establishes true randomness as a consequence of quantum theory. The DIQKD approach thus offers an additional layer of security by requiring no trust in the measurement devices themselves. Given this framework, device-independent protocols make the accurate bounding of entropy under non-asymptotic conditions a central challenge [12]. The need to rigorously quantify the randomness of measurement outcomes—especially in real-world scenarios where noise and finite sample sizes introduce deviations—underscores the importance of non-asymptotic entropy bounds. Such bounds provide critical insights into the level of security achievable in device-independent settings, where only the observed distribution and the validity of quantum theory are assumed [13, 14] (of course in an actual experiment many more assumptions on the devices come into play; we restrict here on the mathematical framework). As research progresses, these entropy bounds are becoming essential for verifying and implementing DIQKD protocols in practical cryptographic systems, thereby expanding the possibilities for secure communication in untrusted environments [15, 16, 17].
In this work, we contribute to the field of device-independent cryptography by introducing a numerical method that provides provable lower bounds on the conditional von Neumann entropy, which quantifies the asymptotic amount of usable randomness extractable from a given experiment. Specifically, we bound the entropy
$H(A|Q_{E})_{\rho_{AQ_{E}}}\coloneqq S(\rho_{AQ_{E}})-S(\rho_{Q_{E}}),$ (2)
where $A$ is a classical random variable held by Alice and $Q_{E}$ is a quantum system held by Eve. Our method enables certification of randomness using only the observed measurement outcomes from Alice and Bob, under the sole assumption that the experiment is governed by quantum theory. The core idea of our method is to transform a novel class of integral representations of the relative entropy [18] into a sequence of non-commutative optimization problems. These problems can be approximated using the Navascués-Pironio-Acín (NPA) hierarchy of semidefinite programs [19, 20]. While the approximation is still computationally demanding for larger systems, our method offers a significant improvement in efficiency compared to previous approaches.
Computing conditional von Neumann entropies in the device independent setting has been an outstanding open problem for quite some time [12]. Only recently the first methods for this problem were developed by Tan et al. [21] and by Brown, Fawzi, and Fawzi [22], shortly after. Both are suffering from serious performance issues when it comes to analyse protocols with more than an minimal amount of inputs and outputs on each side. This limitation is due the fact that both methods require the simultaneous analysis of many non-commutative variables and/or high-degree polynomials. We have the capability to fix this. In its most economical version, our method can give direct estimates on the von Neumann entropy, whilst requiring minimal computational resources. In detail we need the same amount of resources as in the NPA-based computation of min-entropies [23], which is a lower but usually suboptimal bound to the von Neumann entropy. Nevertheless, if more resources can be allocated we can enhance the approximation quality and convergence speed to the optimum.
@@
In general, the demand for such numerical tools has increased recently evermore due to significant progress in entropy accumulation theorems [13, 14], which provide concrete solutions for randomness extraction in finite-size regimes by reducing it to the asymptotic quantity in (2). Thus, after applying an entropy accumulation theorem, the remaining task is to calculate provable lower bounds for the conditional von-Neumann entropy (2) for specific noise models and input-output data scenarios. Furthermore, by confining the analysis to projective operators, our method achieves a high degree of computational efficiency, making it suitable for real-world applications.
This paper is structured as follows: In section Sec. II, we will present the underlying mathematical framework for our method. We will introduce and analyse an integral representation of the underlying optimization problem and introduce approximations, that are linear in the underlying state, see Thm. II.1. This enables us to formulate the asymptotically extractable randomness as a non-commutative polynomial optimization problem, Thm.II.2. The following sections are dedicated to applications. In Sec.III we will revisit the well known 2222-setting. Here, Alice has 2 measurements with 2 outcomes, and Bob has 2 measurements with 2 outcomes. In section Sec. IV we will focus on settings with more measurements and outcomes. In detail, we will compute the asymptotically extractable randomness for settings in which the CGLMP inequality and the I3322 are measured. For these cases, previous numerical methods do not longer perform well [24]. Our method however does, and we demonstrate that direct bounds on the von Neumann entropy certify much higher randomness extraction rates than a comparable ansatz based on the min-entropy. As a last demonstration, presented in section Sec. V we will test our method on real world experimental data. The data comes from a test run of the experiment [7], which firstly demonstrated the feasibility of DI-QKD with distant users. We show that using the statistics of more than 2 measurements can actually increase the amount of certifiable randomness when compared to a simple test of the CHSH inequality [23].
@@ II Bounding the von-Neumann entropy
In the following we denote by $\mathcal{B}(\mathcal{H})$ the set of (bounded) linear operators on a Hilbert space $\mathcal{H}$ and $\mathcal{S}(\mathcal{H})$ the set of quantum states on $\mathcal{H}$, i.e. all positive operators with unit trace. The trace on $\mathcal{B}(\mathcal{H})$ is denoted as $\operatorname{tr}[\cdot]$. The positive semidefinite cone is denoted by $\mathcal{P}(\mathcal{H})$. Moreover, any self adjoint operator $A\in\mathcal{B}(\mathcal{H})$, can be uniquely decomposed as a difference $A=A^{+}-A^{-}$ of Hilbert-Schmidt orthogonal positive operators $A^{+}$ and $A^{-}$. Let $\operatorname{tr}^{+}[A]\coloneqq\operatorname{tr}[A^{+}]$ denote the trace of the positive part of $A$ (similarly $\operatorname{tr}^{-}[A]\coloneqq\operatorname{tr}[A^{-}]=\operatorname{tr}^{+}[-A]$). Note that this is an SDP given by
$\operatorname{tr}^{+}[A]=\sup$ $\operatorname{tr}[PA]$ (3)
$\operatorname{s.t.}$ $0\leq P\leq\mathds{1}.$
@@
by P. E. Frenkel [18]. For two positive operators $\sigma,\rho\in\mathcal{P}(\mathcal{H})$ acting on a Hilbert space $\mathcal{H}$ the r.h.s. of (4) can be expressed by the integration
$D(\rho\|\sigma)=\frac{1}{\ln 2}\big{(}\operatorname{tr}[\rho-\sigma]+\int_{-\infty}^{+\infty}dt\frac{\operatorname{tr}^{-}[(1-t)\rho+t\sigma]}{|t|(t-1)^{2}}\big{)}.$ (5)
It was shown by Jenčová [25], that the formulation (5) also holds for infinite-dimensional yet separable Hilbert spaces. It can be further rewritten in a form that is more convenient for our purposes (see A.1)
@@
$\hskip 28.45274pt+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]\big{)}.$
Here, the boundaries of the integration range $\mu,\lambda\in\mathbb{R}_{\geq 0}$ have to obey the operator valued inequality
@@
In the DI settings we want to analyse, the states $\rho$ and $\sigma$ of interest arise from the protocol under consideration, here a generic choice for valid $\lambda$ and $\mu$ can be given.
For the remainder of this work, we assume tripartite experiments involving Alice, Bob, and Eve, modeled by the axioms of quantum theory as a tripartite state expressed by a positive operator on a Hilbert space with trace equal to $1$,
@@
Importantly, the argument for the device independence specifies nothing concrete about the Hilbert space $\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{QE}$. In the following we follow the example of generating a raw-key in a DIQKD protocol, even though our methods can be applied for a larger class of problems in section III. Afterwards, we briefly discuss some simple generalizations.
In section III we elaborate on this assumption for certain QKD scenarios. In a DIQKD protocol we usually have some measurements $\mathcal{X}$ on $\mathcal{S}(\mathcal{H}_{Q_{A}})$ with finite outcome set $A$ and similarly $\mathcal{Y}$ on $\mathcal{S}(\mathcal{H}_{Q_{B}})$ with finite outcome set $B$ respectively. Measurements, see Fig.1, are generally understood as sets of POVMs $\{M_{a|x}\}_{a\in A}$ for each $x\in\mathcal{X}$ and similarly $\{M_{b|y}\}_{b\in B}$ for each $y\in\mathcal{Y}$. From a more abstract perspective a measurement could also be understood as channel that maps from a commutative algebra into the observable algebra of a quantum system. In the most simplest form of a protocol, the raw-key is now constructed by one specific measurement $\tilde{x}\in\mathcal{X}$. This can be generalized later on [26], but we skip the technical details of this for now. The asymptotically extractable randomness of the fixed measurement $\tilde{x}\in\mathcal{X}$ can be described as [^2] For simplicity we abbreviate $\rho_{AQ_{E}|\tilde{x}}\equiv\rho_{AQ_{E}|X=\tilde{x}}$
@@
Hence, we have to lower bound $D(\rho_{AQ_{E}|\tilde{x}}\|\mathds{1}_{A}\otimes\rho_{Q_{E}})$. Specifically, device independent lower bounds satisfy the following prototypical type of optimization problem
@@
whereby we do not aim to specify the concrete underlying Hilbert space. Moreover, the second line can be understood as information about the joint statistics of Alice and Bob, which is nothing more than (1). A simple but sufficient and common example would be that the $c_{abxyi}\in\mathbb{R}$ coefficients form a Bell inequality (see e.g. [27]). In the following we need to estimate the program (10) with an actually quantifiable program. For this purpose, we use the integral representation (6) and derive an upper bound for the relative entropy (4) with an operator valued optimization problem.
@@
$+\operatorname{tr}[\rho]\ln\lambda-(\lambda-1)\operatorname{tr}[\sigma]).$
@@
The Theorem II.1 provides a sequence of upper bounds that converge to the relative entropy. Specifically, this sequence of optimization problems can be understood as an application of [28]. We emphasize that a finite $\lambda$ is, at least in finite dimensions, equivalent to the statement that the relative entropy is finite. This is a crucial component in these optimization problems, allowing for effective numerical analysis. In the case of conditional von-Neumann entropy, a finite $\lambda$ always exists, as for any bipartite state $\rho_{AB}$ we have the fundamental operator inequality (see, e.g., [29, Appendix A]):
@@
Now, let us examine the specific optimization problem in (10). It can be lower-bounded using Theorem II.1 as an optimization over operators $P_{0},\ldots,P_{r}$. In particular, we observe that the optimal operators $P_{k}$, $0\leq k\leq r$, are projections, as they correspond to projections onto the positive part in the Jordan decomposition of the Hermitian operator $\alpha_{k}\rho+\beta_{k}\sigma$. Furthermore, by substituting the specific operators $\rho=\mathds{1}_{A}\otimes\rho_{Q_{E}}$ and $\sigma=\rho_{AQ_{E}|\tilde{x}}$, we find that $\lambda=1$ and $\mu$ can be set to 0 without loss of generality (see Appendix A for details). Thus, we can rewrite (10) as
@@
An application of Naimark’s dilation theorem additionally allows us to rewrite the state $\rho_{Q_{A}Q_{B}Q_{E}}\in\mathcal{S}(\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}})$ without loss of generality as a pure state $\psi_{Q_{A}Q_{B}Q_{E}}\in\mathcal{H}_{Q_{A}}\otimes\mathcal{H}_{Q_{B}}\otimes\mathcal{H}_{Q_{E}}$, with all measurements also being projective. By carefully examining the classical-quantum structure of $\rho_{AQ_{E}|\tilde{x}}$, the post-measurement state in our specific key basis $\tilde{x}\in\mathcal{X}$ can then be used to reformulate (15) as a non-commutative polynomial optimization in Theorem II.2.
@@
The polynomial optimization problem (16) can be interpreted as an optimization over the algebraic constraints within it. In a finite-dimensional Hilbert space setting, these constraints can be viewed as generating rules of a $\star$-subalgebra of $\mathcal{B}(\mathcal{H})$, where $\mathcal{H}$ is a Hilbert space (in infinite dimensions, caution is required to handle the topological closure of the algebra, though this remains feasible). Assuming that these relations form an algebra $\mathcal{A}$, we employ the Navascues-Pironio-Acin (NPA) hierarchy [19], to compute outer approximations to the optimization problem (16), thus lower bounds.
The approach involves relaxing the positivity requirement for linear, continuous functionals on $\mathcal{A}$. Rather than requiring positivity across the positive cone $\mathcal{A}_{+}$, we restrict it to a sum-of-squares cone $\Sigma_{2}^{(n)}\subset\mathcal{A}_{+}$ at level $n$. Optimization over this relaxed SDP can then be viewed as optimizing the dual cone $(\Sigma_{2}^{(n)})^{\star}$, which is representable in SDP form and provides a convergent hierarchy of outer approximations to (16). For further details, we refer to [19, 30, 20] and Appendix D. In our applications, we use the software package from [31] and code is available here.
0.750.750.80.80.850.8500.50.511CHSH valueconditional entropylower boundanalytic value
0.750.750.80.80.850.8500.50.5111.51.5CHSH valueconditional entropylower boundupper bound [32]
0.80.80.90.91101122white noiseconditional entropylower bound 2322
Figure 2: The plots illustrate examples of lower bounds for conditional entropies across different scenarios. In the leftmost plot, we present a proof-of-principle example for lower bounds on the conditional entropy $H(A|X=0,Q_{E})_{\rho}$, depending on various values of the CHSH violation in a 2222-scenario (where each of two parties can choose between two measurements with two outcomes). This example is particularly significant because it allows for a verification of the general correctness of the method; as shown in [11], there are analytical values available for this scenario, enabling comparisons with computed results to validate accuracy. The middle plot displays the lower bounds on global randomness, specifically the conditional entropy $H(AB|X=0,Y=0,Q_{E})$, again in a 2222-scenario and upper bounds developed by [32]. Here, the dependency is on the CHSH inequality, and this plot serves to examine the impact of the CHSH violation on the joint conditional entropy of outcomes $A$ and $B$ given measurement settings $X=0$ and $Y=0$. Finally, the rightmost plot demonstrates the quantity $H(AB|X=0,Y=0,Q_{E})$ in a 2322-scenario, where each party has two possible outcomes and Bob has three measurement choices. In this case, we condition on the full distribution of a Werner state depending on dephasing noise. All plots can be generated within seconds to a few minutes of computation on a machine equipped with 12 threads and 128GB of RAM.